Hydrogen gas was collected in a tube by water disp
Answer #1

Total pressure of tube is equal to vapor pressure of water and hydrogen pressure

vapor pressure of water + hydrogen pressure    =743mmHg

water vpor pressure = 3Kpa

1KPa = 7.500617mmHg

water vpor pressure = 3*7.500617 = 22.501851 mmHg

vapor pressure of water + hydrogen pressure    =743mmHg

22.501851   + Hydrogen pressure        = 743 mmHg

Hydrogen pressure = 743-22.501851 = 720.498149mmHg

paritial pressure of hydrogen is 720.498149mmHg